# Immunization against harmful fine-tuning attacks

### v1

@@ C.4 Discussion
How can we directly tie these results to the immunization conditions? First, we demonstrated resistance empirically using the stability and harmfulness thresholds estimated from the initial model before the immunization and harmful training. However, these results are not strong enough to show either strong or weak resistance. For strong resistance, we think this will need to be shown theoretically bounds on the performance we can achieve on a proxy harmfulness metric given our immunization method. For weak resistance, this experiment is simply not strong enough, we need to provide much longer training runs to show the maximum step at which we still maintain immunization. Future research should answer the question: What number of training steps or attacks does it take to break immunization? Additionally, in our threat model, the attacker has access to a wide variety of modeling and training tools including hyper-parameter tuning and choice of different types of training objective and architectures like parameter-efficient tuning. A valid and reliable immunization solution for weak resistance will need to show defence across a wide range of different types of training set ups and hyper-parameters.
@@
For the generalization condition, we discussed above some of the desirable empirical settings we need to show generalization. These include: providing a wide range of different harmful domains and attacks that we’d want to illustrate immunization on; the ability to show immunization using as small a number of samples as possible; and extensive analysis of what training-time attacks are not prevented by a given immunization solution. Additionally, while it isn’t the focus of our threat model, we should understand the impact of immunization on test-time attacks as well.
Finally, we did not show trainability in our demonstration. While this isn’t strictly necessary for defences, we believe that LLM developers will want to safely deploy LLMs that can be fine-tuned on harmless datasets and so we challenge the research community to continue to incorporate trainability on their analysis of defences against harmful fine-tuning attack.

### v2

@@ C.4 Discussion
How can we directly tie these results to the immunization conditions? First, we demonstrated resistance empirically using the harmfulness thresholds estimated from the initial model before the immunization and harmful training. However, these results are not strong enough to show either strong or weak resistance. For strong resistance, we think this will need to be shown theoretically. For weak resistance, this experiment is simply not strong enough, we need to provide much longer training runs to show the maximum step at which we still maintain immunization. Future research should answer the question: What number of training steps or attacks does it take to break immunization? Additionally, in our threat model, the attacker has access to a wide variety of modeling and training tools including hyper-parameter tuning and choice of different types of training objective and architectures like parameter-efficient tuning. A valid and reliable immunization solution for weak resistance will need to show defense across a wide range of different types of training set ups and hyper-parameters.
@@
For the generalization condition, we discussed in the main text some of the desirable empirical settings we need to show generalization. These include: providing a wide range of different harmful domains and attacks that we’d want to illustrate immunization on; the ability to show immunization using as small a number of samples as possible; and extensive analysis of what training-time attacks are not prevented by a given immunization solution. Additionally, while it isn’t the focus of our threat model, we should understand the impact of immunization on inference-time attacks as well.
Finally, the adversarial loss method did not show trainability in our demonstration. While this isn’t strictly necessary for defenses, we believe that LLM developers will want to safely deploy LLMs that can be fine-tuned on harmless datasets and so we challenge the research community to continue to incorporate trainability in their analysis of defenses against harmful fine-tuning attack.